# Understanding via Gaze: Gaze-based Task Decomposition for Imitation Learning of Robot Manipulation

### v1

@@ IV-A Robot System
Our group previously developed a dual-arm robotic system designed for imitation learning through teleoperation [8]. A human teleoperator controls the robot remotely by viewing its environment through a head-mounted display (HMD), which simultaneously records the operator’s eye gaze. The system can be configured to use either a physical dual-arm robot (a pair of UR5 arms) or a simulated counterpart, enabling the collection of expert demonstrations in both real and virtual settings [27]. The recorded gaze data are provided as pixel coordinates corresponding to the images displayed on the HMD.
In this study, we evaluated our task decomposition method using an expert demonstration dataset collected by the authors. Each demonstration includes synchronized left and right image streams, along with corresponding gaze data, and all of these were employed in our experiments.
@@ IV-B Task Setup
  • WrapCandy (real robot): In this task, a green candy and a brown handkerchief are arranged on a table with a green tablecloth. The position and orientation of the two objects are randomized on the table. The robot grasps the candy and place it in the middle of the handkerchief. The robot then picks up the (top) left corner of the handkerchief and folds it to wrap the candy.
  • PenInCup (real robot): In this task, a blue marker pen and a red cup are placed on a table; the position and orientation of the two objects is randomized on the table. The robot picks up the pen and puts it in the cup.
  • PileBox (simulator robot): In this task, red and blue boxes are placed on a wooden table; the position and orientation of the two boxes is randomized. The robot grabs the red box and places it on top of the green box.
@@ IV-C Task Decomposition Results
For each task, we performed task decomposition on all demonstrations in the dataset. In doing so, we used the same hyperparameters across all demonstrations and tasks: $w=20,b=256\;(\text{overall image size: }1280\times 720),\theta_{pos}=50,\theta_{feat}=0.05$. Fig. 2 presents an representative example of segmentation results for each task.
@@
  • Each sub-task began when the agent (released the previous object and) started reaching for the next relevant object.
@@
Crucial aspects of object manipulation exist even before actual contact with the object. For example, how the agent reaches for the object significantly influences the success of the manipulation. For this reason, non-contact actions such as reaching should also be included in the segment. However, methods that rely solely on outcome-based cues, such as contact events, cannot detect these reaching actions.
IV-D Uniformity of Task Decomposition
From a practical perspective, it is desirable that all demonstrations within each task exhibit a uniform pattern of task decomposition. Table I shows the number of demonstrations classified as “minority data,” in which the segmentation timings or the number of sub-tasks differed from the most common (majority) decomposition for each task.
In addition, we also evaluated decomposition uniformity under the following ablation settings: (1) without the refinement process (w/o refine), (2) detecting changes based solely on the gaze position $\hat{g}_{t}$ (w/o feat), (3) detecting changes based solely on visual features around the gaze $f_{t}$ (w/o pos), and (4) omitting the refinement process from these two settings (w/o feat, refine and w/o pos, refine).
@@ IV-D Uniformity of Task Decomposition
When the refinement process was applied, we found that all demonstrations across all tasks were segmented in a consistent manner. In contrast, removing the refinement process led to some demonstrations whose decomposition differed from the majority. The number of such minority cases increased noticeably in settings where the gaze position was not used (i.e., w/o pos, refine) compared to those where it was used (i.e., w/o refine and w/o feat, refine). In each scenario, we observed instances of over-segmentation when the gaze shifted slightly during an action, or missed segmentation when the gaze displacement was too small (e.g., when objects were close together). Furthermore, in the condition where only gaze features were employed, change detection was triggered whenever the hand or an object grasped by the hand entered the cropped image region, resulting in over-segmentation.
@@ IV-E Hyperparameter Robustness
Figure 3 shows the number of demonstrations that were correctly segmented when $w$ was set to $\{5,10,15,20,25,30\}$ and $\theta_{pos}$ was set to $\{25,50,75,100,125,150\}$.
When hyperparameters varied extensively, the number of demonstrations that resulted in incorrect decomposition exhibited gradual fluctuations in the absence of refinement processing. In contrast, with refinement processing applied, decomposition was successfully achieved for all demonstrations across the broad parameter range. These findings allow us to conclude that the proposed method possesses the robustness necessary for application to various robotic systems.
In most cases, incorrect decomposition persisted after the refinement process (including instances where the number of successfully segmented demonstrations decreased post-refinement) because decomposition errors resulting from an incorrect number of segments had already become predominant beforehand. Consequently, instead of achieving its intended purpose, the refinement process altered demonstrations that had been correctly segmented initially, leading to an incorrect number of sub-tasks.
Furthermore, setting the window width $w$ to 25 or higher caused the influence of median filtering to become excessively strong. As a result, particularly in the WrapCandy task, it became impossible to detect the required number of segmentation points during the refinement process, regardless of how much the threshold $\theta_{pos}$ was lowered.
@@ V CONCLUSIONS
In this study, we proposed a robust and generalizable method for the autonomous task decomposition of object manipulation behaviors, which serves as a foundation for enabling robots, as autonomous agents, to move beyond rote imitation and to understand appropriately demonstrated object manipulations. This method achieves segmentation solely through detecting the agent’s gaze transitions. Notably, the incorporation of a refinement process allows for consistent task decomposition, which is particularly desirable in applications such as machine learning. Our findings suggest that gaze plays a crucial role in cognition and learning during object manipulation for autonomous agents, including humans. Once a robot acquires gaze control, it can autonomously perform human-like segmentation, which serves as the basis for diverse learning processes, relying solely on the observation of its own gaze transitions.
To achieve greater autonomy in the future, robots must be capable of directly observing and imitating human behaviors rather than relying on teleoperation. In such situations, even if human gaze data were available, using it as a substitute for the robot’s gaze would be inappropriate because humans utilize other sensory modalities, such as tactile information, in addition to visual input for object manipulation. Therefore, breakthroughs in autonomous gaze control are urgently needed today, either for such cases or for cases where demonstration datasets simply lack gaze data from teleoperators, in order to make robots more intelligent.

### v2

@@ IV-A Robot System
We used a dual-arm robotic system that was previously developed for imitation learning through teleoperation [8]. A human teleoperator controlled the robot remotely by viewing its environment through a head-mounted display (HMD), which simultaneously recorded the operator’s eye gaze. The system can be configured to use either a physical dual-arm robot, a pair of UR5 robots from Universal Robots Inc., or a simulated counterpart, enabling the collection of expert demonstrations in both real and virtual settings [30]. The recorded gaze data were provided as pixel coordinates corresponding to the images displayed on the HMD.
In this study, we evaluated our task decomposition method using an expert demonstration dataset collected through this system. Each demonstration included synchronized left and right image streams, along with corresponding gaze data, and all of these were employed in our experiments.
@@ IV-B Task Setup
  • WrapCandy (real robot, 64 demos): In this task, a green candy and a brown handkerchief are arranged on a table with a green tablecloth. The position and orientation of the two objects are randomized on the table. The robot grasps the candy and place it in the middle of the handkerchief. The robot then picks up the (top) left corner of the handkerchief and folds it to wrap the candy.
  • PenInCup (real robot, 125 demos): In this task, a blue marker pen and a red cup are placed on a table; the position and orientation of the two objects is randomized on the table. The robot picks up the pen and puts it in the cup.
  • PileBox (simulated robot, 147 demos): In this task, red and blue boxes are placed on a wooden table; the position and orientation of the two boxes is randomized. The robot grabs the red box and places it on top of the green box.
@@ IV-C Task Decomposition Results
For each task, we performed task decomposition on all demonstrations in the dataset. In doing so, we used the same hyperparameters across all demonstrations and tasks: $w=20,b=256\;(\text{overall image size: }1280\times 720),\theta_{pos}=50,\theta_{feat}=0.05$. Fig. 2 presents a representative example of segmentation results for each task.
@@
  • Each sub-task began when the robot (released the previous object and) started reaching for the next relevant object.
@@
IV-D Consistency of Task Decomposition
In this section, we evaluated the consistency of the resulting task decomposition. Table I shows the number of demonstrations classified as “minority data,” in which the segmentation timings or the number of sub-tasks differed from the most common (majority) decomposition for each task. In addition to the proposed approach, we also evaluated the decomposition consistency under the following ablation settings: (1) without the refinement process (w/o refine), (2) detecting changes based solely on the gaze position $\hat{g}_{t}$ (w/o feat), (3) detecting changes based solely on visual features around the gaze $f_{t}$ (w/o pos), and (4) omitting the refinement process from these two settings (w/o feat, refine and w/o pos, refine).
@@ IV-D Consistency of Task Decomposition
As illustrated in Table I, we found that all demonstrations across all tasks were segmented in a consistent manner when the refinement process was applied. In contrast, removing the refinement process led to some demonstrations whose decomposition differed from the majority. The number of such minority cases increased noticeably in settings where the gaze position was not used (i.e., w/o pos, refine) compared to those where it was used (i.e., w/o refine and w/o feat, refine). In each scenario, we observed instances of over-segmentation when the gaze shifted slightly during an action, or missed segmentation when the gaze displacement was too small (e.g., when objects were close together). Furthermore, in the condition where only gaze features were employed, change detection was triggered whenever the end-effector or a grasped object entered the cropped image region, resulting in over-segmentation.
@@ IV-E Hyperparameter Robustness
Fig. 3 shows the number of demonstrations that were correctly segmented when $w$ was set to $\{5,10,15,20,25,30\}$ and $\theta_{pos}$ was set to $\{25,50,75,100,125,150\}$. When hyperparameters varied extensively, the number of demonstrations that resulted in incorrect decomposition exhibited gradual fluctuations in the absence of refinement processing (Fig. 3). In contrast, with refinement processing applied, decomposition was successfully achieved for all demonstrations across the broad parameter range (Fig. 3). These findings allow us to conclude that the proposed method possesses the robustness necessary for application to various robotic systems.
In cases where incorrect task decomposition persisted after the refinement process, including instances where the number of successfully segmented demonstrations was lower with refinement (Fig. 3) than without it (Fig. 3), decomposition errors resulting from an incorrect number of segments had already become predominant beforehand. Consequently, instead of achieving its intended purpose, the refinement process altered demonstrations that had been correctly segmented initially, leading to an incorrect number of sub-tasks. Furthermore, setting the window width $w$ to 25 or higher caused the influence of median filtering to become excessively strong. As a result, it became impossible to detect the required number of segmentation points during the refinement process, regardless of how much the threshold $\theta_{pos}$ was lowered.
@@ V CONCLUSIONS
In this study, we proposed a practical method that provides a consistent and robust task decomposition for robot manipulation in a recent imitation learning framework. The most important finding is that the desired task decomposition was successfully achieved for all demonstrations by adjusting the threshold in the refinement process. This result suggests that utilizing gaze transitions for task decomposition is fundamentally effective.
However, the proposed method has several limitations. Based on the characteristics of the current imitation learning dataset, we assumed that all demonstrations for a given task consist of the same set of sub-tasks. In contrast, in natural settings, the actions taken to accomplish the same task may consist of a different number of sub-tasks. For example, when pouring water into a cup, one may either grasp the cup before pouring or pour without grasping it. In even more general settings, demonstrations are no longer labeled by task, and a single demonstration may include multiple tasks, as seen in play data [31]. In such cases, our gaze transition detection can be performed, but the refinement process cannot be applied, resulting in several percent incorrect task decomposition.
To achieve greater autonomy in the future, robots must be capable of directly observing and imitating human behaviors rather than relying on teleoperation. In such a scenario, the robot can no longer use the expert’s gaze as supervision. Therefore, breakthroughs in autonomous gaze control are urgently needed so that the robot can determine where to focus during the demonstration in an unsupervised manner.